# latexText page

### v1

@@ 4 Proposed models for estimating dynamic ES contributions
$\bm{w}_{t+1}$ $=\upsilon_{\bm{\theta}}(\bm{w}_{t},\bm{X}_{t}):=\operatorname{ilr}^{-1}\left(\bm{\tau}+\bm{\Phi}\operatorname{ilr}(\bm{w}_{t})+\bm{\Psi}\bm{Y}_{t}\right),\quad t\in\mathbb{T},$ (9)
where $\bm{Y}_{t}$ is a $q$-dimensional random vector representing a vector of covariates, and $\bm{\theta}=(\bm{\tau},\bm{\Phi},\bm{\Psi})$ is a set of parameters on $\Theta$ such that $\bm{\tau}\in\mathbb{R}^{d-1}$, $\bm{\Phi}\in\mathbb{R}^{(d-1)\times(d-1)}$, and $\bm{\Psi}\in\mathbb{R}^{(d-1)\times q}$. The initial weight $\bm{w}_{1}$ can also be regarded as a parameter to be estimated or as a constant determined by some external information. We then estimate parameters in (9) by minimizing the average score of the $d$-tuple of ES contributions in Proposition 1 (S2) with $\phi_{1}=\cdots=\phi_{d}=:\phi$, $\phi(x)=x^{2}$:
$\hat{\bm{\theta}}=\operatorname{argmin}_{\bm{\theta}\in\Theta}\sum_{t\in\mathbb{T}_{\text{in}}}\sum_{j=1}^{d}\left\{x_{j,t}-\upsilon_{\bm{\theta}}(\bm{w}_{t},\bm{x}_{t})_{j}\widehat{ES}_{t}\right\}^{2}\mathbbm{1}{\left\{s_{t}>\widehat{VaR}_{t}\right\}},$ (10)
where $\bm{x}_{t}=(x_{1,t},\dots,x_{d,t})$ and $s_{t}$ are the realizations of $\bm{X}_{t}$ and $S_{t}$, respectively, and $\widehat{VaR}_{t}$ and $\widehat{ES}_{t}$ are the estimates given in the first stage. Once the estimated parameter $\hat{\bm{\theta}}$ is obtained, we forecast the ES contributions at time $n+1$ by ${\widehat{ESC}}_{j,n+1}=\hat{w}_{j,n+1}\widehat{ES}_{n+1}$, $j=1,\dots,d$, where $\hat{\bm{w}}_{1}=\bm{w}_{1}$ and $\hat{\bm{w}}_{t+1}=\upsilon_{\hat{\mathbf{\theta}}}(\hat{\bm{w}}_{t},\bm{X}_{t})$, $t=1,\dots,n$. Note that the estimated weight $\hat{\bm{w}}_{n+1}$ belongs to $\mathcal{S}_{d}$, which ensures the full allocation property $\sum_{j=1}^{d}\widehat{ESC}_{j,t}=\widehat{ES}_{t}$.
@@ 5 Empirical study
Various models in the literature are available to forecast ES contributions. In this section we compare them with our proposed model in Section 4 in the traditional and comparative backtests introduced in Section 2.3 and based on the Murphy diagrams introduced in Section 3. We first describe the setting and models in comparison in Section 5.1. We then show results of the forecast evaluation in Section 5.2. Finally, we discuss forecast accuracy of the compared models in Section 5.3.
@@ 5.1 Setting and model description
We consider a portfolio consisting of the following $d=3$ stock prices with equal weights: Amazon (AMZN), Alphabet Class A (GOOGL), and Telsa (TSLA). From 2010-06-30 to 2023-05-30, each series consists of $n+T=3249$ negative daily log returns multiplied by $100$. We conduct a rolling window analysis with window size $n=2249$ and forecast day-ahead ES contributions for the last $T=1000$ observations. Namely, for $t=n+1,\dots,n+T$, we forecast total VaR, total ES, and ES contributions of $\bm{X}_{t}|\mathcal{G}_{t-1}$ based on the past $n$ observations $\{\bm{X}_{s}\}_{s\in\{t-n,\dots,t-1\}}$. Following the Fundamental Review of the Trading Book (FRTB), we focus on the confidence levels $\alpha=0.975$ (BCBS 2013).
@@
In this analysis we compare $6$ models including the proposed one in Section 4. We briefly introduce these models here and defer a detailed description to Section S2.2. As a simple benckmark, our first model is the historical simulation (HS) model, where at each time $t$ we estimate the risk functional $\bm{\varrho}_{t}=(ESC_{1,t},\dots,ESC_{d,t},VaR_{t},ES_{t})$ nonparametrically based on the past $n$ observations. Second, we consider what we call the bottom-up GARCH (GARCH.BU) model, where estimates of the risk functionals are induced from a copula-GARCH model (Jondeau and Rockinger 2006; Huang et al. 2009) among $\{\bm{X}_{t}\}$. To compare the bottom-up and top-down approaches, we also consider the top-down GARCH (GARCH.TD) model, where we fit $d$ number of bivariate copula-GARCH models on $\{(X_{j,t},S_{t})\}$, $j=1,\dots,d$. Since the above three models do not take the hysteresis effect and time-varying correlations into account, our fourth model is the bivariate hysteretic autoregressive model with GARCH error and dynamic conditional correlations  (HAR.GARCH,  Chen et al. 2019), fitted to $\{(X_{j,t},S_{t})\}$, $j=1,\dots,d$, in the top-down approach. Following (2), total ES is estimated as the sum of estimated ES contributions.
@@
We next include two compositional regression models in our comparison. These models share the estimates $\{\widehat{VaR}_{t},\widehat{ES}_{t}\}$, which are obtained by an AR-GARCH model on $\{S_{t}\}$. Our fifth model is termed the compositional regression model with least square estimation (CR.LSE). In this model we follow Boonen et al. (2019) and first estimate a series of ES contributions under the assumption that each $\bm{X}_{t}|\mathcal{G}_{t-1}$ follows an elliptical distribution. We then transform this series by (8) to obtain the allocation weights $\{\hat{\bm{w}}_{t}\}\subset\mathcal{S}_{d}$. By regarding this set of allocation weights as compositional data, we fit the compositional regression model (9) by the standard least square estimation. Finally, our sixth model is the compositional regression model based on score minimization (CR.OPT), which is the proposed model described in Section 4. We choose the vector of covariates $\bm{Y}_{t}=\left(\operatorname{ilr}\circ\operatorname{C}\left(\max(\bm{X}_{t}S_{t},\bm{\epsilon})\right),\operatorname{ilr}\circ\operatorname{C}\left(-\min(\bm{X}_{t}S_{t},-\bm{\epsilon})\right)\right)$, where $\bm{\epsilon}=(0.01,\dots,0.01)\in\mathbb{R}^{d}$ is a $d$-vector of small numbers, and $\max$ and $\min$ are applied component-wise. The term $\bm{X}_{t}S_{t}=(X_{1,t}S_{t},\dots,X_{d,t}S_{t})$ is motivated by the fact that the allocation weight $\bm{w}_{t}$ equals $\operatorname{C}(\mathbb{E}[\bm{X}_{t}S_{t}|{\cal G}_{t-1}])$ when $\bm{X}_{t}|{\cal G}_{t-1}$ follows an elliptical distribution with zero vector of location parameters. Therefore, we regard the corresponding regression coefficients of $\bm{Y}_{t}$ as the positive and negative effects of the recent covariance term between $\bm{X}_{t}$ and $S_{t}$ to the balance among allocation weights. We introduce $\bm{\epsilon}$ to ensure that all the components of $\max(\bm{X}_{t}S_{t},\bm{\epsilon})$ and $\min(\bm{X}_{t}S_{t},-\bm{\epsilon})$ are non-zero. For the initial allocation weight $\bm{w}_{1}$, we use the compositional data $\hat{\bm{w}}_{1}$ generated in CR.LSE.
@@
In Section S2.3 we display the estimated dynamics of total VaR, total ES, and ES contributions on the out-of-sample period $\mathbb{T}_{\text{out}}$ together with the realized losses. In summary, we observe that HS only captures the average trends, HAR.GARCH tends to estimate ES contributions lower than others, and CR.OPT typically leads to more fluctuated estimates. Compared with these three models, the other three models seem to produce estimates relatively similar with each other.
@@ 5.2 Model evaluation
We conduct the comparative and traditional backtests in the two-step approach. In Section S2.4 we also conduct those in the one-step approach (Fissler and Hoga 2023). For the comparative backtests, we follow Section 2.3 of Nolde and Ziegel (2017) and conduct the Diebold-Mariano (DM)-tests (Diebold and Mariano 1995) for the series of score differences. For an HAC estimator (Andrews 1991) in the test statistic, we choose the Bartlett kernel (Newey and West 1987) with the automatic bandwidth estimator based on AR(1) approximation; see Section 6 of Andrews (1991). In the context of Nolde and Ziegel (2017), we choose CR.OPT for the benchmark model and compare other models as internal models. Consequently, in the three-zone approach of Fissler et al. (2016), the red region indicating superiority of CR.OPT over a compared model is statistically supported, the green region shows that CR.OPT is inferior to the alternative model, and the yellow region means that further investigation is required since there is no statistical evidence on the order between the two models. We fix the significance level to be $0.05$. We repeat this analysis for total VaR, total ES, tuple of ES contributions, and $j$th ES contribution for $j=1,\dots,d$. For the scoring functions, we choose the pinball loss for VaR, AL log score (Taylor 2019) for ES, and square loss for tuple of ES contributions, as well as each of them. Tables 1 and 2 report the average scores and p-values of the one-sided and two-sided DM-tests, based on which we obtain the regions in the three-zone approach.
Table 1: Results of the DM-tests to compare the forecast accuracy of total VaR, total ES, and tuple of ES contributions with CR.OPT as the benchmark model.
[…]
@@
GARCH.BU 48.016 5 0.225 0.888 0.112 yellow
GARCH.TD 46.780 3 0.200 0.900 0.100 yellow
@@
GARCH.BU 393.304 5 0.426 0.787 0.213 yellow
GARCH.TD 391.345 1 0.697 0.348 0.652 yellow
@@
CR.LSE 391.436 2.5 —- —- —- —-
CR.OPT 391.436 2.5 —- —- —- —-
(3) Tuple of ES contributions
HS 189.216 6 0.004 0.998 0.002 red
GARCH.BU 115.213 4 0.062 0.969 0.031 red
GARCH.TD 97.073 3 0.426 0.787 0.213 yellow
HAR.GARCH 124.779 5 0.135 0.932 0.068 yellow
CR.LSE 95.256 2 0.432 0.784 0.216 yellow
CR.OPT 85.428 1 —- —- —- —-
  • ${}^{a}$verage scores are multiplied by $100$.
  • ${}^{b}$anks are based on average scores.
  • ${}^{c}$he p-values are calculated based on the three different null hypotheses, where “$=$ CR.OPT” means that the model is equally accurate as CR.OPT, and $\leq$ ($\geq$) CR.OPT” represents the hypothesis that the model is less (more) accurate than CR.OPT.
  • ${}^{d}$esults of the three-zone approach are presented.
Table 2: Results of the DM-tests to compare the forecast accuracy of $j$th ES contribution, $j\in\{1,\dots,d\}$, with CR.OPT as the benchmark model. See the footnotes in Table 1 for details.

### v2

@@ 4 Proposed models for estimating dynamic ES contributions
$\bm{w}_{t+1}$ $=\upsilon_{\bm{\theta}}(\bm{w}_{s},\bm{X}_{s},s\leq t):=\operatorname{ilr}^{-1}\left(\bm{\tau}+\bm{\Phi}\operatorname{ilr}(\bm{w}_{t})+\bm{\Psi}h(\bm{X}_{s},s\leq t)\right),\quad t\in\mathbb{T},$ (9)
where $h$ is a $\mathbb{R}^{q}$-valued function and $\bm{\theta}=(\bm{\tau},\bm{\Phi},\bm{\Psi})$ is a set of parameters on a parameter space $\Theta$ such that $\bm{\tau}\in\mathbb{R}^{d-1}$, $\bm{\Phi}\in\mathbb{R}^{(d-1)\times(d-1)}$, and $\bm{\Psi}\in\mathbb{R}^{(d-1)\times q}$. For brevity, we only regress $\bm{w}_{t+1}$ by $\bm{w}_{t}$ although one can choose more lagged variables as regressors. The initial weight $\bm{w}_{1}$ can also be regarded as a parameter to be estimated or as an externally given constant. We then estimate parameters in (9) by minimizing the average score of the $d$-tuple of ESCs in Proposition 1 (S2) with $\phi_{1}=\cdots=\phi_{d}=:\phi$, $\phi(x)=x^{2}$:
$\hat{\bm{\theta}}=\operatorname{argmin}_{\bm{\theta}\in\Theta}\sum_{t\in\mathbb{T}_{\text{in}}}\sum_{j=1}^{d}\left\{x_{j,t}-\upsilon_{\bm{\theta}}(\bm{w}_{s},\bm{x}_{s},s\leq t)_{j}\widehat{ES}_{t}\right\}^{2}\mathbbm{1}{\left\{s_{t}>\widehat{VaR}_{t}\right\}},$ (10)
where $\bm{x}_{t}=(x_{1,t},\dots,x_{d,t})^{\top}$ and $s_{t}$ are the realizations of $\bm{X}_{t}$ and $S_{t}$, respectively, and $\widehat{VaR}_{t}$ and $\widehat{ES}_{t}$ are the estimates given in the first stage. Once the estimated parameter $\hat{\bm{\theta}}$ is obtained, we forecast the ESCs at time $n+1$ by ${\widehat{ESC}}_{j,n+1}=\hat{w}_{j,n+1}\widehat{ES}_{n+1}$, $j=1,\dots,d$, where $\hat{\bm{w}}_{1}=\bm{w}_{1}$ and $\hat{\bm{w}}_{t+1}=\upsilon_{\hat{\bm{\theta}}}(\hat{\bm{w}}_{s},\bm{X}_{s},s\leq t)$, $t=1,\dots,n$. Note that the estimated weight $\hat{\bm{w}}_{n+1}$ belongs to $\mathcal{S}_{d}$, which ensures the empirical counterpart of FAP: $\sum_{j=1}^{d}\widehat{ESC}_{j,t}=\widehat{ES}_{t}$.
Remark 3.
If we assume that $ESC_{j,t}<ES_{t}$ instead of $ESC_{j,t}>0$, for all $j=1,\dots,d$ and $t\in\mathbb{T}$, then it holds that $-(d-2)\,ES_{t}<ESC_{j,t}$ by FAP. Therefore, by defining the vector of the normalized weights :
$\tilde{\bm{w}}_{t}=(\tilde{w}_{1,t},\dots,\tilde{w}_{d,t})^{\top}\quad\text{where}\quad\tilde{w}_{j,t}=\frac{ESC_{j,t}+(d-2)\,ES_{t}}{(d-1)^{2}\,ES_{t}},\quad j=1,\dots,d,$
we have that $\tilde{\bm{w}}_{t}\in\left(0,1/(d-1)\right)^{d}$ and that $\sum_{j=1}^{d}\tilde{w}_{j,t}=1$. Since $\tilde{\bm{w}}_{t}$ belongs to a subset of $\mathcal{S}_{d}$, one could naively fit the model (9) with the range constraint $\tilde{\bm{w}}_{t}\in\left(0,1/(d-1)\right)^{d}$ possibly handled through penalization in the loss function in (10). We do not explore this direction further in this study. Indeed, for the portfolio considered in Section 5, we observe that the forecasted ES and ESCs in the whole period are all positive. Hence we adopt the assumption that $ESC_{j,t}>0$ for all $j=1,\dots,d$ and $t\in\mathbb{T}$.
@@ 5 Empirical study
Various models in the literature are available to forecast ESCs. In this section we compare them with our proposed model in Section 4 in the traditional and comparative backtests presented in Section 2.3 and based on the Murphy diagrams introduced in Section 3. We first describe the setting and models in comparison in Section 5.1. We then show results of the forecast evaluation in Section 5.2. Finally, we discuss forecast accuracy of the compared models in Section 5.3.
@@ 5.1 Setting and model description
We consider a portfolio consisting of the following $d=3$ stock prices with equal investment weights (which should not be confused with the ratios of ESCs to the total ES modeled in Section 4): Amazon (AMZN), Alphabet Class A (GOOGL), and Telsa (TSLA). From 2010-06-30 to 2023-05-30, each series consists of $n+T=3249$ negative daily log returns multiplied by $100$. We conduct a rolling window analysis with window size $n=2249$ and forecast day-ahead ESCs for the last $T=1000$ observations. Namely, for $t=n+1,\dots,n+T$, we forecast total VaR, total ES, and ESCs of $\bm{X}_{t}|\mathcal{G}_{t-1}$ based on the past $n$ observations $\{\bm{X}_{s}\}_{s\in\{t-n,\dots,t-1\}}$. Following the Fundamental Review of the Trading Book (FRTB), we focus on the confidence levels $\alpha=0.975$ (BCBS 2013).
@@
In this analysis we compare six models including the proposed one in Section 4. We briefly introduce these models here and defer a detailed description to Section S2.2. As a simple benckmark, our first model is the historical simulation (HS) model, where at each time $t$ we estimate the risk functional $\bm{\varrho}_{t}=(ESC_{1,t},\dots,ESC_{d,t},VaR_{t},ES_{t})$ nonparametrically based on the past $n$ observations. Second, we consider what we call the bottom-up GARCH (GARCH.BU) model, where estimates of the risk functionals are induced from a copula-GARCH model (Jondeau and Rockinger 2006; Huang et al. 2009) among $\{\bm{X}_{t}\}$. We also consider the top-down GARCH (GARCH.TD) model, where we fit $d$ number of bivariate copula-GARCH models on $\{(X_{j,t},S_{t})\}$, $j=1,\dots,d$. Since the above three models do not take the hysteresis effect and time-varying correlations into account, our fourth model is the bivariate hysteretic autoregressive model with GARCH error and dynamic conditional correlations  (HAR.GARCH,  Chen et al. 2019), fitted to $\{(X_{j,t},S_{t})\}$, $j=1,\dots,d$, in the top-down approach. Following (2), total ES is estimated as the sum of estimated ESCs in this model.
@@
We next include two compositional regression models in our comparison. These models share the estimates $\{\widehat{VaR}_{t},\widehat{ES}_{t}\}$, which are obtained by an AR-GARCH model on $\{S_{t}\}$ used in the above GARCH.TD model for the purpose of comparing these models. Our fifth model is termed the compositional regression model with least square estimation (CR.LSE). In this model we follow Boonen et al. (2019) and first estimate a series of ESCs under the assumption that each $\bm{X}_{t}|\mathcal{G}_{t-1}$ follows an elliptical distribution. We then transform this series by (8) to obtain the allocation weights $\{\hat{\bm{w}}_{t}\}\subset\mathcal{S}_{d}$. By regarding this set of allocation weights as compositional data, we fit the compositional regression model (9) by the standard least square estimation. Finally, our sixth model is the compositional regression model based on score optimization (CR.OPT), which is the proposed model described in Section 4. To preserve interpretability of the model, we choose $\operatorname{ilr}$-transformed variables for $h(\bm{X}_{t})$. More specifically, we choose $h(\bm{X}_{s},s\leq t)=(h^{+}(\bm{X}_{s},s\leq t),h^{-}(\bm{X}_{s},s\leq t))^{\top}$ with
$h^{+}(\bm{X}_{s},s\leq t)$ $=\operatorname{ilr}\circ\operatorname{C}\left(\max\left(\frac{1}{t_{0}}\sum_{s=0}^{t_{0}-1}\bm{X}_{t-s}S_{t-s},\bm{\epsilon}\right)\right),$
$h^{-}(\bm{X}_{s},s\leq t)$ $=\operatorname{ilr}\circ\operatorname{C}\left(-\min\left(\frac{1}{t_{0}}\sum_{s=0}^{t_{0}-1}\bm{X}_{t-s}S_{t-s},-\bm{\epsilon}\right)\right),$
where $t_{0}=7$, $\bm{\epsilon}=(0.01,\dots,0.01)\in\mathbb{R}^{d}$ is a $d$-vector of small numbers, and $\max$ and $\min$ are applied component-wise. The moving average term $(1/t_{0})\,\sum_{s=0}^{t_{0}-1}\bm{X}_{t-s}S_{t-s}$ is motivated by the fact that the allocation weight $\bm{w}_{t}$ equals $\operatorname{C}(\mathbb{E}[\bm{X}_{t}S_{t}|{\cal G}_{t-1}])$ when $\bm{X}_{t}|{\cal G}_{t-1}$ follows an elliptical distribution with zero vector of location parameters. Therefore, we regard the corresponding regression coefficients of $h(\bm{X}_{t})$ as the positive and negative effects of the local covariance between $\bm{X}$ and $S$, calculated over the past $t_{0}$ days, to the balance among allocation weights. In particular, for the reduced model:
$\operatorname{ilr}(\bm{w}_{t+1})=\operatorname{ilr}(\tilde{\bm{\tau}})+\phi\operatorname{ilr}(\bm{w}_{t})+\psi^{+}\operatorname{ilr}(h^{+}(\bm{X}_{s},s\leq t))+\psi^{-}\operatorname{ilr}(h^{-}(\bm{X}_{s},s\leq t)),$
where $\tilde{\bm{\tau}}=\operatorname{ilr}^{-1}(\bm{\tau})$ and $\phi,\psi^{+},\psi^{-}\in\mathbb{R}$, the vector of weights $\bm{w}_{t+1}$ is given by the power-perturbation combination of $\bm{w}_{t}$, $h^{+}(\bm{X}_{s},s\leq t)$ and $h^{-}(\bm{X}_{s},s\leq t)$, the counterpart of linear combination in a simplicial geometry; see, for example, Aitchison and J. Egozcue (2005). We extend this model by allowing mutual effects between $\operatorname{ilr}_{1}$ and $\operatorname{ilr}_{2}$. We introduce $\bm{\epsilon}$ to ensure that all the components of $\max(\bm{X}_{t}S_{t},\bm{\epsilon})$ and $\min(\bm{X}_{t}S_{t},-\bm{\epsilon})$ are non-zero. More sophisticated treatment with zeros can be possible; see Chapter 4 of Pawlowsky-Glahn and Buccianti (2011) and references therein. For the initial allocation weight $\bm{w}_{1}$, we use the compositional data $\hat{\bm{w}}_{1}$ generated in CR.LSE.
@@
In Figures 1 and 2 we display the estimated dynamics of total VaR, total ES, and ESCs on the out-of-sample period $\mathbb{T}_{\text{out}}$ together with the realized losses. Note that the forecasted VaR, ES and ESCs in the whole period are all greater than 00. To summarize our observations on the models, we find that HS only captures the average trends, HAR.GARCH tends to estimate ESCs lower than others, and CR.OPT typically leads to more fluctuated estimates. Compared with these three models, the other three models seem to produce estimates relatively similar with each other.
Figure 1: Time series plots of total VaR and total ES with confidence level $0.975$ estimated by the six models. The black lines represent the time series of the total loss of the portfolio, and the losses exceeding the estimated total VaRs used in CR.OPT are marked in blue.
Figure 2: Time series plots of ESCs with confidence level $0.975$ estimated by the six models. The black lines represent the time series of the corresponding marginal losses, and the losses exceeding the estimated total VaRs used in CR.OPT are marked in blue.
In Table 1 we provide summary statistics of the estimated parameters in CR.LSE and CR.OPT over the backtesting period. According to the table, the parameters estimated in CR.OPT are overall more volatile over time compared with those in CR.LSE. For both of CR.LSE and CR.OPT, the diagonal elements of $\bm{\Phi}$ tend to take large values than the off-diagonal elements. In addition, the effect of $h(\bm{X}_{s},s\leq t)$ parametrized by $\bm{\Psi}$ is estimated to be constantly small in CR.LSE whereas it is sometimes significant in CR.OPT.
Table 1: Summary statistics of the parameters of the model $\operatorname{ilr}(\bm{w}_{t+1})=\bm{\tau}+\bm{\Phi}\operatorname{ilr}(\bm{w}_{t})+\bm{\Psi}^{+}\operatorname{ilr}(h^{+}(\bm{X}_{s},s\leq t))+\bm{\Psi}^{-}\operatorname{ilr}(h^{-}(\bm{X}_{s},s\leq t))$ over the backtesting period, where $\bm{\tau}=(\tau_{1},\tau_{2})^{\top}$, $\bm{\Phi}=(\phi_{ij})_{2\times 2}$, $\bm{\Psi}^{+}=(\psi^{+}_{ij})_{2\times 2}$ and $\bm{\Psi}^{-}=(\psi^{-}_{ij})_{2\times 2}$.
CR.LSE CR.OPT
Mean Median SD 1st Qu. 3rd Qu. Mean Median SD 1st Qu. 3rd Qu.
$\tau_{1}$0.003 -0.003 0.003 -0.005 -0.001 0.045 0.037 0.045 0.018 0.063
$\tau_{2}$0.043 -0.044 0.009 -0.050 -0.035 -0.041 -0.039 0.027 -0.056 -0.025
$\phi_{11}$ 0.793 0.790 0.021 0.779 0.801 0.839 0.876 0.119 0.800 0.912
$\phi_{21}$ 0.027 0.031 0.010 0.017 0.035 0.084 0.052 0.135 -0.005 0.146
$\phi_{12}$0.010 -0.007 0.009 -0.010 -0.004 -0.071 -0.073 0.094 -0.130 -0.033
$\phi_{22}$ 0.901 0.898 0.018 0.889 0.903 0.877 0.891 0.072 0.839 0.928
$\psi^{+}_{11}$ 0.053 0.053 0.004 0.049 0.057 -0.051 -0.036 0.070 -0.073 -0.015
$\psi^{+}_{21}$ 0.011 0.011 0.002 0.009 0.013 0.058 0.060 0.039 0.038 0.080
$\psi^{+}_{12}$ 0.006 0.006 0.002 0.004 0.007 -0.014 -0.012 0.036 -0.039 0.010
$\psi^{+}_{22}$ 0.035 0.038 0.006 0.036 0.039 0.082 0.083 0.036 0.053 0.108
$\psi^{-}_{11}$ 0.053 0.052 0.005 0.049 0.056 -0.002 0.019 0.135 -0.056 0.082
$\psi^{-}_{21}$ 0.005 0.005 0.002 0.003 0.007 0.056 0.051 0.104 -0.006 0.107
$\psi^{-}_{12}$ 0.006 0.007 0.003 0.003 0.009 0.051 0.020 0.167 -0.041 0.116
$\psi^{-}_{22}$ 0.043 0.046 0.008 0.043 0.048 0.111 0.101 0.085 0.057 0.157
@@ 5.2 Model evaluation
We conduct the comparative and traditional backtests in the two-step approach, where we handle hypotheses on VaR (and ES) and those on ESCs separately; see Section 5.4 in (Fissler and Hoga 2024). In Section S2.3 we also conduct those in the alternative one-step approach (Fissler and Hoga 2024). For the comparative backtests, we follow Section 2.3 of Nolde and Ziegel (2017) and conduct the Diebold-Mariano (DM)-tests (Diebold and Mariano 1995) for the series of score differences. For an HAC estimator (Andrews 1991) in the test statistic, we choose the Bartlett kernel (Newey and West 1987) with the automatic bandwidth estimator based on AR(1) approximation; see Section 6 of Andrews (1991). In the context of Nolde and Ziegel (2017), we choose CR.OPT for the benchmark model and compare other models as internal models. Consequently, in the three-zone approach of Fissler et al. (2016), the red region indicates that the superiority of CR.OPT over a compared model is statistically supported, the green region shows that CR.OPT is inferior to the alternative model, and the yellow region means that further investigation is required since there is no statistical evidence on the order between the two models. We fix the significance level to be $0.05$. We repeat this analysis for total VaR, total ES, tuple of ESCs, and $j$th ESC for $j=1,\dots,d$, although, in terms of multi-objective elicitability, the results on ESCs are not meaningful if forecast accuracy is ranked in total VaRs. For the scoring functions, we choose the pinball loss for VaR, AL log score (Taylor 2019) for ES, and squared loss for tuple of ESCs, as well as each of them. Tables 2 and 3 report the average scores and p-values of the one-sided and two-sided DM-tests, based on which we obtain the regions in the three-zone approach.
Table 2: Results of the DM-tests to compare the forecast accuracy of total VaR, total ES, and tuple of ESCs with CR.OPT as the benchmark model.
Average scorea Rankb p-valuec Regiond
@@
GARCH.BU 47.903 5 0.254 0.873 0.127 yellow
GARCH.TD 46.721 1 —- —- —- —-
@@
GARCH.BU 393.015 5 0.498 0.751 0.249 yellow
GARCH.TD 391.436 1 —- —- —- yellow
@@
CR.LSE 391.436 1 —- —- —- —-
CR.OPT 391.436 1 —- —- —- —-
(3) Tuple of ESCs
HS 189.216 6 0.003 0.998 0.001 (red)
GARCH.BU 118.196 4 0.021 0.990 0.010 red
GARCH.TD 94.699 2 0.255 0.872 0.128 yellow
HAR.GARCH 124.779 5 0.202 0.899 0.101 yellow
CR.LSE 95.256 3 0.086 0.957 0.043 red
CR.OPT 85.839 1 —- —- —- —-
  • aAverage scores are multiplied by $100$.
  • bRanks are based on average scores.
  • cThe p-values are calculated based on the three different null hypotheses, where “$=$ CR.OPT” means that the model is equally accurate as CR.OPT, and $\leq$ ($\geq$) CR.OPT” represents the hypothesis that the model is less (more) accurate than CR.OPT.
  • dResults of the three-zone approach are presented. For (3), the result of HS is enclosed by parentheses since the order on total VaR is already supported in (1).
Table 3: Results of the DM-tests to compare the forecast accuracy of $j$th ESC, $j\in\{1,\dots,d\}$, with CR.OPT as the benchmark model. See the footnotes in Table 2 for details.